OpenStudy (anonymous):

Solve this ordinary differential equation by using the method of variation of parameters y''(x) + y(x) = sec x.

OpenStudy (accessdenied):

Our first step should be to solve the homogeneous DE associated with this DE: \(y' ' + y = 0\) The independent solutions of this equation will be useful in the variatioin of parameters solution method.

OpenStudy (accessdenied):

The method of finding the particular solution is using this structure. \(Y_p = u_1 y_1 + u_2 y_2 \) For linearly independent solutions to the homogeneous DE \(y_1\) and \(y_2\) (we find from above) We could work out the solution with some manipulations, or there are also the formulas for \(u_1\) and \(u_2\). \(y" + p(x) y' + q(x) y = f(x) \) \(\to \quad \displaystyle u_1 = - \int \frac{y_2 f(x)}{W(x)} \; \text{d}x \) \( \quad \quad \ \displaystyle u_2 = \int \frac{y_1 f(x)}{W(x)} \; \text{d}x \) W(x) the wronskian of y_1 and y_2.

OpenStudy (accessdenied):

Then the general solution to the main problem is: \( y = c_1 y_1 + c_2 y_2 + Y_p \) Just piecing together those smaller things we found before.

OpenStudy (accessdenied):

Would this all make sense? Is there anything I should go over in more detail? :)

OpenStudy (anonymous):

thanks dear....

OpenStudy (accessdenied):

You're welcome!